OpenStudy (anonymous):

If customer A has an appointment 2 2/3 hours before customer B who has an appointment 5/6 of an hour before customer C. How much later is customer C‘s appointment than customer A’s?

OpenStudy (anonymous):

3 1/2

OpenStudy (anonymous):

wow you good thx

OpenStudy (across):

\[2\frac{2}{3}=\frac{8}{3}\]\[\frac{8}{3}+\frac{5}{6}\]\[\frac{16}{6}+\frac{5}{6}\]\[\frac{21}{6}=3\frac{3}{6}=3\frac{1}{2}\]superjew917 is correct.

OpenStudy (anonymous):

thx